7. Let’s assume Best Buy is a variable growth stock with an expected annual growth rate in dividends of 10% in year 2, 14% in years 3 & 4, and 13.5% in year 5 and a constant annual growth rate of 8% after year 5. What is your valuation of Best Buy’s stock today using the variable growth dividend discount model? Would you recommend buying BEST BUY’s stock today and why? Use current price in #6 to help aid your recommendation. (5 pts)

Below is the current price from #6Best Buy Activision Stock $55.40 $68.40 Current Price Estimated Price 1 year from now $75.89 $57.52 $2.00 $0.37 Annual Expect
Answer #1

Dividend in year 2 (D2) = D1 * 10% + D1 = 2*10%+10 = $2.2

Price of the stock in 1 year from now (P1) = $75.89

P1 = D2/(r-g2), g2 is the growth rate of dividend in year 2

75.89 = 2.2/(r-10%)

r = Cost of Equity = 12.9%

Now we calculate all the dividends,

D1 is given = $2

D2 = $2.2

D3 = 2.2*(1+14%) = $2.51

D4 = 2.51*(1+14%) = $2.86

D5 = 2.86*(1+13.5%) = $3.25

D6 = 3.25*(1+8%) = $3.51

P5 = D6/(r-gc), gc = constant growth rate after year 5

P5 = 3.51/(12.9%-8%) = $71.63

Stock Valuation today = D1/(1+r) + D2/((1+r)^2) + D3/((1+r)^3) + D4/((1+r)^4) + (D5+P5)/((1+r)^5) = $47.82

Since the current stock price of $68.4 is more than the Stock Valuation of $47.82 the stock is overvalued currently and it is not recommended to buy it.